OpenStudy (anonymous):

An object is launched upward at 64 ft/sec from a platform that is 80 feet high. What is the objects maximum height if the equation of height (h) in terms of time (t) of the object is given by h(t) = -16t^2 + 64t + 80?

OpenStudy (campbell_st):

you need to find the time that the max height occurs. have you done calculus...?

OpenStudy (anonymous):

no, just algebra 2

OpenStudy (campbell_st):

ok... that helps.... there are a few methods I'd recommend you find the line of symmetry for a parabola \[at^2 +bt + c = 0\] the line of symmetry is \[t = \frac{-b}{2a}\] you have a = -16 and b = 64 can you substitute them to find t, the time the max height occurs..?

OpenStudy (campbell_st):

an alternate method is to divide every term by -16 and then put the parabola into the vertex form y = (x - h)^2 + k

OpenStudy (anonymous):

what's x, h and k?

OpenStudy (campbell_st):

(h, k) is the vertex, it can be used as an alternate method. my advice is to use the 1st method I posted. Which is using the line of symmetry.

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

but that leaves me with -2??? @campbell_st

OpenStudy (anonymous):

how is that correct? did I do something wrong? @Campbell_st

OpenStudy (campbell_st):

it leaves you with t = 64/(-2x-16) so at t = 2 the pumpkin reaches its maximum height. so then find h(2) = -16x(2)^2 + 64 x 2 + 80 this will give the max height.